# Generalized and Unified Equivalences between Hardness and Pseudoentropy

### v1

@@ 1.4.3 Exponential Lower Bound
In our Theorem 1.10, we make a computational efficiency assumption about the convex functions $\varphi\in\Phi$ corresponding to the entropy notions considered in the theorem: we assume that the subgradient function $\nabla\varphi$ can be approximated by a circuit of size at most $T_{\Phi}$ for every $\varphi\in\Phi$. Correspondingly, the circuit complexity of the function $s$ guaranteed by the theorem has circuit complexity depending polynomially on $T_{\Phi}$, in addition to the polynomial dependency on $T,L$ and $1/\varepsilon$.
@@
For every sufficiently large positive integer $n$, choosing $L=n$, there exist a distribution $\mu$ over $\{0,1\}^{n}$, a function $g^{*}:\{0,1\}^{n}\to\Delta_{L}$, an $O(n)$-sized circuit $f:\{0,1\}^{n}\to[-1,1]^{L}$, and a convex function $\varphi:\Delta_{L}\to\mathbb{R}$ with bounded subgradient $\nabla\varphi(v)\in[-1,1]^{L}$ for every $v\in\Delta_{L}$ that satisfy the following property. Let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function that is ($(\{f\},0.05)$-indistinguishable from $g^{*}$ (i.e. $s\in\mathsf{Ind}(g^{*};\{f\},0.05)$) and satisfies the following inequality:
@@
We would like to remark that the condition (7) is a very mild assumption. In particular, by the nonnegativity of Bregman divergences, (7) is a necessary condition for the following guarantee of our main theorem (Theorem 1.10):
@@
Therefore, Theorem 1.12 implies an exponential lower bound on the circuit complexity of every function $s\in\mathsf{Ind}(g^{*};\{f\},0.05)$ that satisfies the guarantee (4) of Theorem 1.10, even when $\Phi$ only contains a single function $\varphi$ and there is only a single, extremely simple distinguisher $f$ (the identity function).
There is no contradiction between the $\exp(\Omega(L))$ lower bound (Theorem 1.12) and the $\mathsf{poly}(L)$ upper bound (Theorem 1.10) on the circuit complexity of $s$ because the upper bound assumes that the convex function $\varphi$ has low complexity (specifically, its subgradient can be approximated by a circuit of size $T_{\Phi}$). Thus, Theorem 1.12 shows that some assumptions on the complexity of $\varphi$ are necessary for Theorem 1.10 to hold, even when $\Phi$ only contains a single convex function $\varphi$.
Our proof of Theorem 1.12 uses coding-theoretic ideas (specifically, the existence of a sufficiently large design) as well as a probabilistic counting argument. See Section 6 for the formal proof.
As a corollary of Theorem 1.12, we give an exponential lower bound on the circuit complexity for achieving multiaccuracy (i.e. indistinguishability) plus weight-restricted calibration, even for a single, extremely simple distinguisher $f$ (the identity function) and a single weight function $\nabla\varphi:\Delta_{L}\to[-1,1]^{L}$. Consequently, this lower bound also holds for stronger notions such as calibrated multiaccuracy and multicalibration. See Section 1.5 for a more detailed discussion about these notions from the algorithmic fairness literature (multiaccuracy, weight-restricted calibration, calibrated multiaccuracy, and multicalibration).
@@
For every sufficiently large positive integer $n$, choosing $L=n$, the same distribution $\mu$, function $g^{*}$, distinguisher $f$ and convex function $\varphi$ from Theorem 1.12 have the following property. Let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function satisfying the following two conditions:
@@ 1.5 Proof Idea: Regularity Lemma with Weight-Restricted Calibration
We now give a high-level description of our proof idea for the first part of Theorem 1.10, i.e. the forward direction (4). Our proof of the reverse direction (5) follows the same ideas as in prior work [31, 34].
Our idea is very simple in hindsight. As mentioned earlier, a key difference between our guarantee (4) from previous ones (e.g. (1)) is that the same function $s$ appears on both sides of the inequality, playing the roles of the simulator and the approximator simultaneously. This makes it natural to calculate the diference between the two sides. Specifically, using the definitions of $\mathsf{H}_{\varphi}$ and $\mathsf{D}_{\varphi}$ in Definitions 1.7 and 1.8, we obtain the following identity, which is the central technical insight of our work:
@@
such that $s$ is efficiently computable and indistinguishable from $g^{*}$. In Theorem 1.10, we assume that $\nabla\varphi$ is bounded in $[-1,1]^{L}$ and can be approximated within $\ell_{\infty}$ error $\varepsilon/4$ in non-uniform time $T_{\Phi}$. Thus for every $\varphi\in\Phi$, there exists $r_{\varphi}:\Delta_{L}\to[-1,1]$ computable in non-uniform time $T_{\Phi}$ such that $\|r_{\varphi}-\nabla\varphi\|_{\infty}\leq\varepsilon/4$. This gives a sufficient condition for (10):
[…]
As we will explain shortly, the condition (11) is termed weight-restricted calibration in the recent literature on algorithmic fairness. On the other hand, the requirement that $s$ is indistinguishable from $g^{*}$ (Definition 1.6) amounts to
[…]
@@
The first half of Theorem 1.10 requires us to find a function $s\in\mathsf{Ind}(g^{*};T,\varepsilon)\cap\mathsf{Time}(T^{\prime})$ satisfying (4). Based on the our discussion above, the requirement $s\in\mathsf{Ind}(g^{*};T,\varepsilon)$ is equivalent to (12), and a sufficient condition for (4) is given by (11). Thus our goal becomes finding $s\in\mathsf{Time}(T^{\prime})$ satisfying both (11) and (12).
The structure of (11) and (12) points us to previous ideas about regularity lemmas. Specifically, the Complexity-Theoretic Regularity Lemma of Trevisan, Tulsiani, and Vadhan [29] and its extension to $L>2$ in the Leakage Simulation Lemma of Jetchev and Pietrzak [22], allows us to construct $s\in\mathsf{Time}(T^{\prime})$ that satisfies (12) alone. This result can be proved using a boosting algorithm. We start with a simple $s:\{0,1\}^{n}\to\Delta_{L}$ (e.g. $s(x)$ is the uniform distribution over $[L]$ for every $x\in\{0,1\}^{n}$). Whenever $s$ does not satisfy (12), there exists a (low-complexity) $f\in\mathcal{F}$ that witnesses this violation, which we use to update $s$ to reduce a nonnegative potential function by a significant amount. By the nonnegativity of the potential function, there can only be a bounded number of updates, after which $s$ has to satisfy (12). Since each $f\in\mathcal{F}$ has non-uniform complexity at most $T$, each update increases the complexity of $s$ only by a bounded amount, and since the number of total updates is also bounded, we can ensure that the final $s$ belongs to $\mathsf{Time}(T^{\prime})$. Following the work of Vadhan and Zheng [32] which proves a uniform version of the regularity lemma, we apply the classic Multiplicative Weights algorithm (a special case of Mirror Descent) to extend the boosting idea from [29] to general $L$ (see Sections 2.2 and 2.3).
@@
A key insight from the recent algorithmic fairness literature about multicalibration [15, 12] is that the circularity in (11) does not incur any new major technical challenge: the exact same boosting idea from the Regularity Lemma allows us to construct $s\in\mathsf{Time}(T^{\prime})$ satisfying both (11) and (12), which is exactly what we need to establish Theorem 1.10. We view condition (11) as an instantiation of weight-restricted calibration, as introduced by Gopalan, Kim, Singhal, and Zhao [12]. [^4] They called it weighted calibration (see also [10]). We modify the terminology to more clearly convey that it is a weaker condition than calibration. This is a relaxed version of the following condition called calibration:
[…]
@@
[…]
where each $r:\Delta_{L}\to[-1,1]^{L}$ is termed a weight function [12]. The weight-restricted calibration condition (11) is a relaxation of (full) calibration (14) in that it only considers a restricted family of weight functions $\{r_{\varphi}\}_{\varphi\in\Phi}$ instead of all weight functions (and it replaces “$\approx 0$” with “$\leq\varepsilon$”).
@@
The combination of multiaccuracy (12) and (full) calibration (13) (or (14)) is termed calibrated multiaccuracy in [9]. That work uses calibrated multiaccuracy to achieve omniprediction more efficiently than the multicalibration-based approach of [11] (we discuss multicalibration in the next paragraph). The notion of multiaccuracy (12) plus weight-restricted calibration (11) (as we use in our work) is weaker than calibrated multiaccuracy and suffices to ensure omniprediction in many cases, leading to further improved sample complexity and computational efficiency even when $L=2$ [20, 26]. Independent work [2, 6] demonstrates the power of (variants of) calibrated multiaccuracy for other cryptographic, complexity-theoretic, and learning-theoretic applications (as we discussed in Section 1.2).
@@
[…]
A contribution of our work is to show that multicalibration is an overkill for pseudoentropy characterizations and leads to the unnecessary doubly exponential dependence on $\ell$ in Theorem 1.5 (our Corollary 1.13 shows that this doubly exponential dependence cannot be avoided for multicalibration and even for the weaker notion of calibrated multiaccuracy). Based on our observation (9), the much weaker condition of multiaccuracy (12) plus weight-restricted calibration (11) is sufficient even for more general entropy notions $\mathsf{H}_{\varphi}$, giving the improved polynomial dependence on $L=2^{\ell}$ in Theorem 1.10.
@@ 1.6 Examples beyond Shannon and Min Entropy
Our pseudoentropy characterization (Theorem 1.10) holds for a general family $\Phi$ of entropy notions. Here we give two examples of entropy notions beyond Shannon and min entropy.
@@
where the first term $\Pr[g^{*}(X)=g(X)]$ is the prediction accuracy of $g$ w.r.t. $g^{*}$ (the same as in Theorem 1.2):
@@
where $X\sim\mu$, and conditioned on $X$, we independently draw $Y\sim g^{*}(X),\hat{Y}\sim g(X)$. The second term on the right-hand side of (16) can be viewed as a regularizer that encourages $g$ to have lower collision probability (and thus higher entropy $\mathsf{H}_{\varphi}$). Our characterization Theorem 1.10 implies that the best that an efficient $g$ can do for this regularized optimization task is equivalent to the conditional pseudo-collision-probability of $g^{*}(X)$ given $X$ (i.e., the pseudo / computational analogue of $\mathsf{CP}(g^{*}(X)|X)$).
@@
$\varphi(v)=\|v\|_{2}=\sqrt{\sum_{i}v_{i}^{2}},\quad\mathsf{D}_{\varphi}(u\|v)=\frac{\|u\|_{2}\|v\|_{2}-\langle u,v\rangle}{\|v\|_{2}}\$
@@ 2 Preliminaries
For simplicity, we consider general and abstract vectors $g,h,f\in\mathbb{R}^{L}$ in this section. Later we will instantiate the vector $g$ here as the function value $g(x)\in\Delta_{L}\subseteq\mathbb{R}^{L}$ (as in Definition 1.6) assigned to a specific input $x\in\{0,1\}^{n}$, and similarly instantiate the vector $f$ here as the function value $f(x)\in[-1,1]^{L}$ of a distinguisher $f:\{0,1\}^{n}\to[-1,1]^{L}$ on a specific input $x\in\{0,1\}^{n}$.
@@ 2.1 Convex Analysis Basics
By the boundedness of $\mathcal{G}$ and the convexity of $\varphi$, the supremum above is always well-defined (i.e., it is neither +$\infty$+ ∞ nor $\infty$- ∞). It is easy to verify that $\psi$ is convex and closed.555 As standard terminology in convex analysis, we say $\psi$ is closed if its epigraph $\{(h,y)\in\mathbb{R}^{L}\times\mathbb{R}:y\geq\psi(h)\}$ is closed (see e.g. [27]). The Fenchel-Young divergence $\Gamma_{\varphi,\psi}:\mathcal{G}\times\mathbb{R}^{L}\to\mathbb{R}$ is defined as follows:

### v2

@@ 1.4.3 Exponential Lower Bound
In our Theorem˜1.10, we make a computational efficiency assumption about the convex functions $\varphi\in\Phi$ corresponding to the entropy notions considered in the theorem: we assume that the subgradient function $\nabla\varphi$ can be approximated by a circuit of size at most $T_{\Phi}$ for every $\varphi\in\Phi$. Correspondingly, the circuit complexity of the function $s$ guaranteed by the theorem depends polynomially on $T_{\Phi}$, in addition to the polynomial dependency on $T,L$ and $1/\varepsilon$.
@@
For every sufficiently large positive integer $n$, choosing $L=n$, there exist a distribution $\mu$ over $\{0,1\}^{n}$, a function $g^{*}:\{0,1\}^{n}\to\Delta_{L}$, an $O(n)$-sized circuit $f:\{0,1\}^{n}\to[-1,1]^{L}$, and a convex function $\varphi:\Delta_{L}\to\mathbb{R}$ with bounded subgradient $\nabla\varphi(v)\in[-1,1]^{L}$ for every $v\in\Delta_{L}$ that satisfy the following property. Let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function that is $(\{f\},0.05)$-indistinguishable from $g^{*}$ (i.e. $s\in\mathsf{Ind}(g^{*};\{f\},0.05)$) and satisfies the following inequality:
@@
We would like to remark that the condition (7) is a very mild assumption. In particular, by the nonnegativity of Bregman divergences, (7) is a necessary condition for the following guarantee of our main theorem (Theorem˜1.10):
@@
Therefore, Theorem˜1.12 implies an exponential lower bound on the circuit complexity of every function $s\in\mathsf{Ind}(g^{*};\{f\},0.05)$ that satisfies the guarantee (4) of Theorem˜1.10, even when $\Phi$ only contains a single function $\varphi$ and there is only a single, extremely simple distinguisher $f$ (the identity function).
There is no contradiction between the $\exp(\Omega(L))$ lower bound (Theorem˜1.12) and the $\mathsf{poly}(L)$ upper bound (Theorem˜1.10) on the circuit complexity of $s$ because the upper bound assumes that the convex function $\varphi$ has low complexity (specifically, its subgradient can be approximated by a circuit of size $T_{\Phi}$). Thus, Theorem˜1.12 shows that some assumptions on the complexity of $\varphi$ are necessary for Theorem˜1.10 to hold, even when $\Phi$ only contains a single convex function $\varphi$.
Our proof of Theorem˜1.12 uses coding-theoretic ideas (specifically, the existence of a sufficiently large design) as well as a probabilistic counting argument. See Section˜6 for the formal proof.
As a corollary of Theorem˜1.12, we give an exponential lower bound on the circuit complexity for achieving multiaccuracy (i.e. indistinguishability) plus weight-restricted calibration, even for a single, extremely simple distinguisher $f$ (the identity function) and a single weight function $\nabla\varphi:\Delta_{L}\to[-1,1]^{L}$. Consequently, this lower bound also holds for stronger notions such as calibrated multiaccuracy and multicalibration. See Section˜1.5 for a more detailed discussion about these notions from the algorithmic fairness literature (multiaccuracy, weight-restricted calibration, calibrated multiaccuracy, and multicalibration).
@@
For every sufficiently large positive integer $n$, choosing $L=n$, the same distribution $\mu$, function $g^{*}$, distinguisher $f$ and convex function $\varphi$ from Theorem˜1.12 have the following property. Let $s:\{0,1\}^{n}\to\Delta_{L}$ be an arbitrary function satisfying the following two conditions:
@@ 1.5 Proof Idea: Regularity Lemma with Weight-Restricted Calibration
We now give a high-level description of our proof idea for the first part of Theorem˜1.10, i.e. the forward direction (4). Our proof of the reverse direction (5) follows the same ideas as in prior work [31, 34].
Our idea is very simple in hindsight. As mentioned earlier, a key difference between our guarantee (4) from previous ones (e.g. (1)) is that the same function $s$ appears on both sides of the inequality, playing the roles of the simulator and the approximator simultaneously. This makes it natural to calculate the diference between the two sides. Specifically, using the definitions of $\mathsf{H}_{\varphi}$ and $\mathsf{D}_{\varphi}$ in Definitions˜1.7 and 1.8, we obtain the following identity, which is the central technical insight of our work:
@@
such that $s$ is efficiently computable and indistinguishable from $g^{*}$. In Theorem˜1.10, we assume that $\nabla\varphi$ is bounded in $1,1]^{L}$ and can be approximated within $\ell_{\infty}$ error $\varepsilon/4$ in non-uniform time $T_{\Phi}$. Thus for every $\varphi\in\Phi$, there exists $r_{\varphi}:\Delta_{L}\to[-1,1]$ computable in non-uniform time $T_{\Phi}$ such that $\|r_{\varphi}-\nabla\varphi\|_{\infty}\leq\varepsilon/4$. This gives a sufficient condition for (10):
$\langle s-g^{*},r_{\varphi}\circ s\rangle\leq\varepsilon/2\quad\text{for every $\varphi\in\Phi$}.\quad\textbf{(Weight-restricted calibration)}$ (11)
As we will explain shortly, the condition (11) is termed weight-restricted calibration in the recent literature on algorithmic fairness. On the other hand, the requirement that $s$ is indistinguishable from $g^{*}$ (Definition˜1.6) amounts to
$|\langle s-g^{*},f\rangle|\leq\varepsilon\quad\text{for every $f\in\mathcal{F}$},\quad\textbf{(Indistinguishability, a.k.a.\ multiaccuracy)}$ (12)
@@
The first half of Theorem˜1.10 requires us to find a function $s\in\mathsf{Ind}(g^{*};T,\varepsilon)\cap\mathsf{Time}(T^{\prime})$ satisfying (4). Based on the our discussion above, the requirement $s\in\mathsf{Ind}(g^{*};T,\varepsilon)$ is equivalent to (12), and a sufficient condition for (4) is given by (11). Thus our goal becomes finding $s\in\mathsf{Time}(T^{\prime})$ satisfying both (11) and (12).
The structure of (11) and (12) points us to previous ideas about regularity lemmas. Specifically, the Complexity-Theoretic Regularity Lemma of Trevisan, Tulsiani, and Vadhan [29] and its extension to $L>2$ in the Leakage Simulation Lemma of Jetchev and Pietrzak [22], allows us to construct $s\in\mathsf{Time}(T^{\prime})$ that satisfies (12) alone. This result can be proved using a boosting algorithm. We start with a simple $s:\{0,1\}^{n}\to\Delta_{L}$ (e.g. $s(x)$ is the uniform distribution over $[L]$ for every $x\in\{0,1\}^{n}$). Whenever $s$ does not satisfy (12), there exists a (low-complexity) $f\in\mathcal{F}$ that witnesses this violation, which we use to update $s$ to reduce a nonnegative potential function by a significant amount. By the nonnegativity of the potential function, there can only be a bounded number of updates, after which $s$ has to satisfy (12). Since each $f\in\mathcal{F}$ has non-uniform complexity at most $T$, each update increases the complexity of $s$ only by a bounded amount, and since the number of total updates is also bounded, we can ensure that the final $s$ belongs to $\mathsf{Time}(T^{\prime})$. Following the work of Vadhan and Zheng [32] which proves a uniform version of the regularity lemma, we apply the classic Multiplicative Weights algorithm (a special case of Mirror Descent) to extend the boosting idea from [29] to general $L$ (see Sections˜2.2 and 2.3).
@@
A key insight from the recent algorithmic fairness literature about multicalibration [15, 12] is that the circularity in (11) does not incur any new major technical challenge: the exact same boosting idea from the Regularity Lemma allows us to construct $s\in\mathsf{Time}(T^{\prime})$ satisfying both (11) and (12), which is exactly what we need to establish Theorem˜1.10. We view condition (11) as an instantiation of weight-restricted calibration, as introduced by Gopalan, Kim, Singhal, and Zhao [12]. [^4] They called it weighted calibration (see also [10]). We modify the terminology to more clearly convey that it is a weaker condition than calibration. This is a relaxed version of the following condition called calibration:
$\mathbb{E}_{x\sim\mu}[s(x)-g^{*}(x)|s(x)]\approx\mathbf{0},\quad\textbf{(Calibration)}$ (13)
@@
$\langle s-g^{*},r\circ s\rangle\approx 0\quad\text{for \emph{all} $r:\Delta_{L}\to[-1,1]^{L}$},\quad\textbf{(Equivalent def.\ of calibration)}$ (14)
where each $r:\Delta_{L}\to[-1,1]^{L}$ is termed a weight function [12]. The weight-restricted calibration condition (11) is a relaxation of (full) calibration (14) in that it only considers a restricted family of weight functions $\{r_{\varphi}\}_{\varphi\in\Phi}$ instead of all weight functions (and it replaces $\approx 0$” with $\leq\varepsilon$”).
@@
The combination of multiaccuracy (12) and (full) calibration (13) (or (14)) is termed calibrated multiaccuracy in [9]. That work uses calibrated multiaccuracy to achieve omniprediction more efficiently than the multicalibration-based approach of [11] (we discuss multicalibration in the next paragraph). The notion of multiaccuracy (12) plus weight-restricted calibration (11) (as we use in our work) is weaker than calibrated multiaccuracy and suffices to ensure omniprediction in many cases, leading to further improved sample complexity and computational efficiency even when $L=2$ [20, 26]. Independent work [2, 6] demonstrates the power of (variants of) calibrated multiaccuracy for other cryptographic, complexity-theoretic, and learning-theoretic applications (as we discussed in Section˜1.2).
@@
$\mathbb{E}_{x\sim\mu}[\langle s(x)-g^{*}(x),f(x)\rangle|s(x)]\approx 0\quad\text{for every $f\in\mathcal{F}$}.\quad\textbf{(Multicalibration)}$
A contribution of our work is to show that multicalibration is an overkill for pseudoentropy characterizations and leads to the unnecessary doubly exponential dependence on $\ell$ in Theorem˜1.5 (our Corollary˜1.13 shows that this doubly exponential dependence cannot be avoided for multicalibration and even for the weaker notion of calibrated multiaccuracy). Based on our observation (9), the much weaker condition of multiaccuracy (12) plus weight-restricted calibration (11) is sufficient even for more general entropy notions $\mathsf{H}_{\varphi}$, giving the improved polynomial dependence on $L=2^{\ell}$ in Theorem˜1.10.
@@ 1.6 Examples beyond Shannon and Min Entropy
Our pseudoentropy characterization (Theorem˜1.10) holds for a general family $\Phi$ of entropy notions. Here we give two examples of entropy notions beyond Shannon and min entropy.
@@
where the first term $\Pr[g^{*}(X)=g(X)]$ is the prediction accuracy of $g$ w.r.t. $g^{*}$ (the same as in Theorem˜1.2):
@@
where $X\sim\mu$, and conditioned on $X$, we independently draw $Y\sim g^{*}(X),\hat{Y}\sim g(X)$. The second term on the right-hand side of (16) can be viewed as a regularizer that encourages $g$ to have lower collision probability (and thus higher entropy $\mathsf{H}_{\varphi}$). Our characterization Theorem˜1.10 implies that the best that an efficient $g$ can do for this regularized optimization task is equivalent to the conditional pseudo-collision-probability of $g^{*}(X)$ given $X$ (i.e., the pseudo / computational analogue of $\mathsf{CP}(g^{*}(X)|X)$).
@@
$\varphi(v)=\|v\|_{2}=\sqrt{\sum_{i}v_{i}^{2}},\quad\mathsf{D}_{\varphi}(u\|v)=\frac{\|u\|_{2}\|v\|_{2}-\langle u,v\rangle}{\|v\|_{2}}\\$
@@ 2 Preliminaries
For simplicity, we consider general and abstract vectors $g,h,f\in\mathbb{R}^{L}$ in this section. Later we will instantiate the vector $g$ here as the function value $g(x)\in\Delta_{L}\subseteq\mathbb{R}^{L}$ (as in Definition˜1.6) assigned to a specific input $x\in\{0,1\}^{n}$, and similarly instantiate the vector $f$ here as the function value $f(x)\in[-1,1]^{L}$ of a distinguisher $f:\{0,1\}^{n}\to[-1,1]^{L}$ on a specific input $x\in\{0,1\}^{n}$.
@@ 2.1 Convex Analysis Basics
By the boundedness of $\mathcal{G}$ and the convexity of $\varphi$, the supremum above is always well-defined (i.e., it is neither $\infty$ nor -$\infty$). It is easy to verify that $\psi$ is convex and closed.555 As standard terminology in convex analysis, we say $\psi$ is closed if its epigraph $\{(h,y)\in\mathbb{R}^{L}\times\mathbb{R}:y\geq\psi(h)\}$ is closed (see e.g. [27]). The Fenchel-Young divergence $\Gamma_{\varphi,\psi}:\mathcal{G}\times\mathbb{R}^{L}\to\mathbb{R}$ is defined as follows: